A rectangle has an area of 98 cm2. It length is double its width. How wide is it?
Answer by lwsshak3(11628)

A rectangle has an area of 98 cm2. It length is double its width. How wide is it?
let W=width
2W=length
area=length*width
98=2W*W=2W^2
W^2=98/2=49
W=√49=7
How wide is rectangle=7 cm